I need help. Here is the problem I have to solve.
Gina is paid 1.5 times normal hourly pay for each she works over 40 hours in a week. Last week she worked 48 hours and earned $634.40. What is her normal hourly rate?
Please help me.

Answer by checkley71(8403)

SHE WORKS 40 HOURS FOR X AND 8 HOURS FOR 1.5X SO THE EQUATION IS 40X+8*1.5X=634.40 OR 40X+12X=634.40 OR 52X=634.40 OR X=634.6/52 OR X=$12.20 WHICH IS HER HOURLY RATE.